if p,q,r,s are any four consecutive terms of an A.P. Show that
p²-3q²+3r²-s² = 0

Answer by Edwin McCravy(20056)

If d is the common difference in the A.P., then

q = p+d,  r = p+2d,  s = p+3d

p²-3q²+3r²-s² = 

p² - 3(q²-r²) - s² =

(p²-s²) - 3(q²-r²) =

Factor both parentheses as the difference of squares:

[p-s][p+s] - 3[q-r][q+r] =

Substitute q = p+d,  r = p+2d,  s = p+3d

[p-(p+3d)][p+(p+3d)] - 3[(p+d)-(p+2d)][(p+d)+(p+2d)] =

[p-p-3d][p+p+3d] - 3[p+d-p-2d][p+d+p+2d] =

[-3d][2p+3d] - 3[-d][2p+3d] =

-3d[2p+3d] + 3d[2p+3d] =

-6dp -9d² + 6dp + 9d² =
   
          0

Edwin